Please calculate the amount of a compound (A) that remains in 10 mL of an aqueous phase after two extractions with toluene (10 mL each) if KD = 2 (KD = partition ratio toluene / water). Show your calculation(s).
Answer #1

To determine the amount of compound A remaining in the aqueous phase after two extractions with toluene, we can use the partition coefficient (KDK_D) and the volumes of the aqueous and organic phases.

Given:

  • Partition coefficient, KD=2K_D = 2 (toluene/water)

  • Initial volume of aqueous phase, Vaq=10mLV_{\text{aq}} = 10 \, \text{mL}

  • Volume of toluene used in each extraction, Vtol=10mLV_{\text{tol}} = 10 \, \text{mL}

  • Number of extractions, n=2n = 2

Assumption:

  • Let the initial amount of compound A in the aqueous phase be C0C_0.

Calculation:

The fraction of compound A remaining in the aqueous phase after each extraction can be determined using the formula:

Fraction remaining after one extraction=11+KD×VtolVaq\text{Fraction remaining after one extraction} = \frac{1}{1 + K_D \times \frac{V_{\text{tol}}}{V_{\text{aq}}}}

Substituting the given values:

Fraction remaining after one extraction=11+2×1010=11+2=13\text{Fraction remaining after one extraction} = \frac{1}{1 + 2 \times \frac{10}{10}} = \frac{1}{1 + 2} = \frac{1}{3}

After one extraction, 13\frac{1}{3} of the compound remains in the aqueous phase.

After two extractions, the fraction remaining is:

(13)2=19\left( \frac{1}{3} \right)^2 = \frac{1}{9}

Therefore, after two extractions, 19\frac{1}{9} of the initial amount of compound A remains in the aqueous phase.

Conclusion:

After two extractions with 10 mL of toluene each, approximately 11.11% of the initial amount of compound A remains in the 10 mL aqueous phase.


answered by: Shital Garg